Circular No. 8774

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2006 S10 AND C/2006 S11 (SOHO) Two additional Kreutz sungrazing comets have been found on SOHO

website images (cf. IAUC 8772). K. Battams notes that C/2006 S10 was small and stellar in appearance (mag ∼ 6) in C3 coronagraph images, while in C2 images it appeared very small, faint, and diffuse. C/2006 S11 was small with a hint of tail and peaked in brightness at mag ∼ 5.5 in C3 images, but it was very diffuse and elongated in C2 images.

Comet  2006 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2006 S10 Sept. 18.763 11^{h}23.^{m}9 + 0^{o}38′ C3/2 RM 2006-U63
C/2006 S11  18.821 11 24.9 + 0 43 C3/2 RM 2006-U63

COMET C/2006 V1 (CATALINA) Regarding an apparently asteroidal object found the previous night

in the course of the Catalina Sky Survey, E. J. Christensen reports that four stacked 60-s CCD exposures obtained in poor seeing with the 1.5-m reflector at Mount Lemmon on Nov. 12.5 UT show an asymmetric 8′′ coma, slightly extended toward p.a. 310^{o}. Following posting by the Minor Planet Center on its ‘NEOCP’ webpage, J. Young writes that his CCD images taken with the Table Mountain 0.61-m reflector on Nov. 15.43–15.48 UT show a round 6′′-diameter coma with little or no central condensation and a suspected very weak, narrow, straight tail in p.a. ∼ 300^{o}. R. A. Kowalski reports that CCD images obtained with the Mount Lemmon 1.5-m reflector on Nov. 15.5 in poor seeing show a slightly condensed nucleus with a coma diameter of ≈ 11′′.

2006 UT  α_{2000}  δ_{2000}  Mag.
Nov. 11.49270  10^{h}55^{m}05.^{s}44  +10^{o}48′36′′.7  18.5

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-V63.

T = 2007 Nov. 24.545 TT  ω = 244.^{o}398
Ω = 336.244 }2000.0
q = 3.06870 AU  i = 30.017
2006 November 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green